Using the Bisection method, find an approximate root of the equation sin(x)=1/x that lies between x=1 and x=1.5 (in radians). Compute upto 5 iterations. Determine the approximate error in each iteration. Give the final answer in a tabular form.
Answer #1

An f(x) = x sinx-1 f(1) = sin() - 1 --O. 15849 20 f(1.5) = 1.5 sin (1.5) -1 = 0.49 625 > 0 The root lies between 1 and 1.5. Ff(x2) = f(1.125) = 1.125 Sin (1.125)-1 = 0. 01509 so Root lies between s and 2 = 1.125. So approximate error Fa - Present apApproximate eroor is Eaz = f(x4) - f(az) = -0.02836 - (-0.8-768) = 0.04344. .. Fifth approximate to the groot is 2 = 1.09375